HETAV PATEL

Grade 11,

while calculating volume strenth of H 2 o 2=normality*5.6 from where this 5.6 has come . .

while calculating volume strenth of H2o2=normality*5.6 from where this 5.6 has come  . . 

Grade:12th pass

1 Answers

Bhavya
8 years ago
We know that Normality(N) = weight(w) / equivalent weight(E)
N*E = w
Now we know that 2 moles of H2O2 gives 1 mole of O2
i.e 68g H2O2 gives 22.4 l O2
‘w’ gH2O2 gives (22.4/68)*w
=(22.4/68)*N*E
= (22.4/68)*17*N
= 5.6*N